OpenStudy (anonymous):

Three vertices of parallelogram DFGH are D(-9, 4) , F(-1, 5) , and G(2, 0). Find the coordinates of vertex H. choices: a. (-6, -1) b. (-9, 0) c. (-2, -3) d. (-1, 5)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

H=(-6,-1) I will explain to you why

OpenStudy (anonymous):

thank you so much~

OpenStudy (anonymous):

Do you want more explanation?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

D(-9, 4) , F(-1, 5) , and G(2, 0). H(-6,1) Find the slope of FG which is equal to the slope FH find the equation of the line FH you find it equal to \[ y=-\frac{5 x}{3}-11 \] Do the same for the line GH to find \[ y=\frac{x}{8}-\frac{1}{4} \] The intersection of the two lines is H. Solve \[ \frac{5 x}{3}-11= \frac{x}{8}-\frac{1}{4} \] you find x=-6 Hence y=-1

OpenStudy (anonymous):

Discard H(-6,1) in the first line of the above post

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

ya thankyou

OpenStudy (anonymous):

Did you understand the method I desribed above